# Multi-Stage Balanced Distillation: Addressing Long-Tail Challenges in Sequence-Level Knowledge Distillation

### v1

@@ Abstract
Large language models (LLMs) have significantly advanced various natural language processing tasks, but deploying them remains computationally expensive. Knowledge distillation (KD) is a promising solution, enabling the transfer of capabilities from larger teacher LLMs to more compact student models. Particularly, sequence-level KD, which distills rationale-based reasoning processes instead of merely final outcomes, shows great potential in enhancing students’ reasoning capabilities. However, current methods struggle with sequence-level KD under long-tailed data distributions, adversely affecting generalization on sparsely represented domains. We introduce the Multi-Stage Balanced Distillation (BalDistill) framework, which iteratively balances training data within a fixed computational budget. By dynamically selecting representative head domain examples and synthesizing tail domain examples, BalDistill achieves state-of-the-art performance across diverse long-tailed datasets, enhancing both the efficiency and efficacy of the distilled models.
@@ 2 Related Work
Knowledge Distillation uses the outputs of a larger LLM (Teacher), such as ChatGPT OpenAI (2023), to train a smaller model (Student), such as LLaMa-7B Touvron et al. (2023). For details of knowledge distillation (KD) of large language models, we refer to the survey for more details  Xu et al. (2024). In this work, we focus on KD with black-box teacher models. There are two lines of work with respect to knowledge distillation. The first is to ask teacher models to generate the final answers and to fine-tune on the final answers  Zhou et al. (2023); Schick and Schütze (2021). Another line of work asks teacher models to generate rationales at the reasoning process and fine-tunes student models on the rationales in the sequence level to improve their reasoning ability Ho et al. (2022); Shridhar et al. (2022); Hsieh et al. (2023), which proves to be more effective. In this work, we mainly discuss using a teacher model to generate rationales and improve the student’s reasoning ability on a long-tailed dataset. Despite the progress of KD in the LLM era, existing works fail to establish a pipeline to gain knowledge from long-tailed datasets with the sequence-level KD, as few rationale examples are provided for tail knowledge.
Long-Tail Learning focuses on long-tail distributed data and has been an emerging topic of interest in the NLP community  Liu et al. (2019); Wang et al. (2017); Godbole and Jia (2022); Dai et al. (2023); Zhang et al. (2022); Liu et al. (2024c). Approaches to solving the long-tail problem include rebalancing, information augmentation, and module improvement  Zhang et al. (2021); He et al. (2021); Zhu et al. (2023); Cui et al. (2021). Despite the importance of long-tail learning, studies have shown that LLMs struggle to learn long-tail knowledge  Kandpal et al. (2023); Sun et al. (2023a). In this work, we propose to improve LLMs’ ability to learn long-tail knowledge via multi-stage distillation over balanced datasets.
Active Learning aims to reduce labeling effort by selecting only the most useful examples. Traditional active learning can be categorized into uncertainty-based methods  Prabhu et al. (2019); Margatina et al. (2021); Wang et al. (2023) and diversity-based methods  Ru et al. (2020); Ash et al. (2019). In the LLM era, active learning has been used to reduce human annotation costs by (1) strategically selecting the most informative examples for human feedback or annotation  Margatina et al. (2023); Osband et al. (2022); Wang et al. (2020) and (2) integrating language models as annotators within an active learning framework without human supervision  Xiao et al. (2023); Rouzegar and Makrehchi (2024); Zhang et al. (2023); Li et al. (2024c). In this work, we propose to solve the long-tail problem in the student LLM by actively distilling knowledge from a black-box teacher LLM to meet the budget requirement.
@@ 3.2 Overall Approach
After selecting and/or generating these samples, we query the teacher model to provide detailed rationales for examples in the training batch. These annotated examples are then used to fine-tune the student model $\mathcal{M}_{s}$ in preparation for the next stage. Detailed descriptions of these components, along with the algorithms outlining this procedure, are presented in Algorithm 1 in Appendix B.
@@ 3.5 Student Active Selection
At last, we rank the inputs by their IFD scores, selecting those with the highest values to include in the bath, as specified by the balancing policy.
@@ 3.6 Reasoning Generation and Fine-tuning
After gathering the required samples and their associated rationales in the training batch, we integrate this batch with the annotated data accumulated from previous stages. This approach ensures that our student model is exposed to a diverse and comprehensive dataset, which helps mitigate the risk of overfitting — a common challenge in machine learning models as identified in prior studies Dor et al. (2020). To facilitate this, we reinitialize and fine-tune the student model on the compiled rationale sequences from scratch at each stage.
@@ 4 Experiment
To verify the effectiveness of our framework on various reasoning tasks, we evaluate our method on five long-tailed datasets, following previous work Yu et al. (2023); Dai et al. (2023); Huang et al. (2021). These include text classification: R52 and Reuters Hayes and Weinstein (1990), question answering: AbstractiveQA and Multiple-choiceQA Dai et al. (2023) and arithmetic: MATH Hendrycks et al. (2021). For text classification datasets, we treat the label of inputs as the domain; for other datasets, the domain information of inputs is annotated as metadata from the data provider. The detailed construction process and domain information for these datasets can be found in Appendix A. We also show two example distributions of the datasets in Figure 5 in Appendix A. For each dataset, we prepare two budget settings for the experiment. In Table 1, we present the budget number, the test number, the domain number, and the evaluation metric of all five datasets.
@@
Since we are dealing with long-tailed imbalanced data, for each dataset, we choose to use both the micro- and macro-averages to evaluate the method robustness Henning et al. (2022). For the classification datasets (R52 and Reuters), we report micro-F1 and macro-F1, where micro-F1 is a global average F1 score and macro-F1 is computed by taking the unweighted mean of all the per-class F1 scores Harbecke et al. (2022). For other datasets, we also report the micro-/macro-F1 for AbstractiveQA datasets and micro-/macro-accuracy for Math and Multi-choiceQA datasets. Note that the F1 score for the AbstractQA is the word-level F1 score between the token list of ground truth answer and the generated answer, different from the F1 for the classification task.
@@
We experiment with two variants of our proposed method with different balancing policies, as discussed in Section 3: In our first framework BalDistill (N), we use naive balancing policy, and for second framework BalDistill (A), we leverage adaptive balancing. We compare our framework with multiple baseline methods: (1) Zero-shot CoT. We directly prompt the student model to infer on the test data Kojima et al. (2022). (2) Random Finetune. We randomly collect samples from the training data until the budget constraint is met and finetune student models on the final ground-truth labels Radford et al. (2019). (3) Random Finetune-CoT. We randomly collect and use CoT rationales from the teacher model for student fine-tuning Ho et al. (2022); Yao et al. (2022); He et al. (2023). (4) Duplicate Finetune-CoT. We construct the training data with a naive balancing policy, and for tail domains, we duplicate the inputs to satisfy the policy requirement.
@@ 5.1 Comparison with Baseline Methods
To perform a detailed analysis of our framework, we visualize the F1 or accuracy score for each domain of the BalDistill (N) method and two baseline methods (Random Finetune CoT and Duplicate Finetune CoT) in Figure 2, with the x-axis representing the proportion of each domain in the dataset in descending order. From Figure 2, our proposed method can achieve comparable results in the head domains (left side of the figure) but substantially outperform the baseline methods in the tail domains (right side of the figure). This observation verifies our expectation in Section 3, where the balancing policy increases performance in the tail domain, and the active learning part improves the data efficiency to compensate for data loss in the head domain. Note that for Math dataset, BalDistill can only achieve comparable results with the baseline methods on the last two tail domains (precalculus and probability), and we conjecture that the high difficulty in these two domains prevents the teacher from composing high-quality synthetic data.
@@ 5.2 Ablation Study
After showing the superiority of our overall framework, our next step is to verify the effectiveness of each component in the proposed method. We compare our framework with the ablated methods: (1) Balance Finetune CoT. We adopt a naive balancing policy to construct the training set and query the teacher model to compose inputs in the tail domains. (2) Active Finetune CoT. We only keep the active learning component but remove the data augmentation part. The experiment setting is similar to the setup in Section 5.1, and we present the performance of each method with two budget settings in Table 3.
@@ 5.3.2 Sensitivity Analysis
We next investigate how the choice of stage number: $K$ will influence the performance of our framework. We experiment with the same setup as in Section 5.1 but with varying stage numbers among $\{3,5,8\}$. We visualize the results (macro-averages) of BalDistill (A) and the baseline method Random Finetune CoT in Figure 4 and 6 in Appendix.
@@
Figure 4: Influence of stage number choices on BalDistill across two datasets. Our proposed method consistently obtains better results than the random fine-tune baseline method with varying stage numbers.
@@ 6 Conclusions
7 Limitations
@@
Another future direction for our paper is to explore the application of knowledge distillation in Large Vision-Language Models (LVLMs) Liu et al. (2024b, a); Bai et al. (2023); Sun et al. (2023b); Zhou et al. (2024a); Wang et al. (2024a); Lin et al. (2024). In this paper, we have focused on experiments related to knowledge distillation in Large Language Models (LLMs). In future work, we aim to use knowledge distillation to further address the issue of hallucination Liu et al. (2023); Cui et al. (2023); Wang et al. (2024b) in small LVLMs such as LLaVA-7b Liu et al. (2024b) and VILA-7b Lin et al. (2024).
@@ Appendix A Dataset Construction
Algorithm 1 Multi Stage Balanced Distillations
2:Output: The fine-tuned student model $M_{s}^{K}$
3:for each stage $k=0,\dots,k-1$ do
[…]
5:     for each domain tail domain $j$ do
6:         Add remaining $x_{j}$ from $D$ to $T$
7:         $\hat{x_{j}}=M_{t}(P_{c},j)$
8:         Add synthetic $\hat{x_{j}}$ to $T$
9:     for each domain head domain $h$ do
10:         Collect all $x_{h}$ from $D$
11:         $x_{h}=M_{s}^{k-1}(x_{h},h)$
12:         Add selected $x_{h}$ to $T$
13:     Use $M_{t}$ to annotate $x$ in $T$ w/o rationales
14:     $M_{s}^{k}=\text{Fine-tune}(M_{s},T)$
@@ Appendix B Implementation Details
(a) MultiChoiceQA
(b) Math
Figure 6: Influence of stage number choices on BalDistill across other two datasets.
@@
We use the zero-shot prompts for the teacher to give the rationales and the few-shot ICL to generate the synthetic tail samples. The prompts are shown in Tables 6, 7 and 8. We call the gpt-4 function from OpenAI to obtain teacher responses.

### v2

@@ Abstract
Large language models (LLMs) have significantly advanced various natural language processing tasks, but deploying them remains computationally expensive. Knowledge distillation (KD) is a promising solution, enabling the transfer of capabilities from larger teacher LLMs to more compact student models. Particularly, sequence-level KD, which distills rationale-based reasoning processes instead of merely final outcomes, shows great potential in enhancing students’ reasoning capabilities. However, current methods struggle with sequence-level KD under long-tailed data distributions, adversely affecting generalization on sparsely represented domains. We introduce the Multi-Stage Balanced Distillation (BalDistill) framework, which iteratively balances training data within a fixed computational budget. By dynamically selecting representative head domain examples and synthesizing tail domain examples, BalDistill achieves state-of-the-art performance across diverse long-tailed datasets, enhancing both the efficiency and efficacy of the distilled models.  [^1] Our code and data are available at https://github.com/Tonyzhou98/long_tail_kd
@@ 2 Related Work
Knowledge Distillation uses the outputs of a larger LLM (Teacher), such as ChatGPT OpenAI (2023), to train a smaller model (Student), such as LLaMa-7B Touvron et al. (2023). For details of knowledge distillation (KD) of large language models, we refer to the survey for more details  Xu et al. (2024b). In this work, we focus on KD with black-box teacher models. There are two lines of work with respect to knowledge distillation. The first is to ask teacher models to generate the final answers and to fine-tune on the final answers  Zhou et al. (2023); Schick and Schütze (2021). Another line of work asks teacher models to generate rationales at the reasoning process and fine-tunes student models on the rationales in the sequence level to improve their reasoning ability Ho et al. (2022); Shridhar et al. (2022); Hsieh et al. (2023), which proves to be more effective. In this work, we mainly discuss using a teacher model to generate rationales and improve the student’s reasoning ability on a long-tailed dataset. Despite the progress of KD in the LLM era, existing works fail to establish a pipeline to gain knowledge from long-tailed datasets with the sequence-level KD, as few rationale examples are provided for tail knowledge.
Long-Tail Learning focuses on long-tail distributed data and has been an emerging topic of interest in the NLP community  Liu et al. (2019); Wang et al. (2017); Godbole and Jia (2022); Dai et al. (2023); Zhang et al. (2022); Liu et al. (2024d); Mondal et al. (2024). Approaches to solving the long-tail problem include rebalancing, information augmentation, and module improvement  Zhang et al. (2021); He et al. (2021); Zhu et al. (2023); Cui et al. (2021); Xu et al. (2024a). Despite the importance of long-tail learning, studies have shown that LLMs struggle to learn long-tail knowledge  Kandpal et al. (2023); Sun et al. (2023a). In this work, we propose to improve LLMs’ ability to learn long-tail knowledge via multi-stage distillation over balanced datasets.
Active Learning aims to reduce labeling effort by selecting only the most useful examples. Traditional active learning can be categorized into uncertainty-based methods  Prabhu et al. (2019); Margatina et al. (2021); Wang et al. (2023) and diversity-based methods  Ru et al. (2020); Ash et al. (2019). In the LLM era, active learning has been used to reduce human annotation costs by (1) strategically selecting the most informative examples for human feedback or annotation  Margatina et al. (2023); Osband et al. (2022); Wang et al. (2020) and (2) integrating language models as annotators within an active learning framework without human supervision  Xiao et al. (2023); Rouzegar and Makrehchi (2024); Zhang et al. (2023); Li et al. (2024c); Liu et al. (2024f). In this work, we propose to solve the long-tail problem in the student LLM by actively distilling knowledge from a black-box teacher LLM to meet the budget requirement.
@@ 3.2 Overall Approach
After selecting and/or generating these samples, we query the teacher model to provide detailed rationales for examples in the training batch. These annotated examples are then used to fine-tune the student model $\mathcal{M}_{s}$ in preparation for the next stage. Detailed descriptions of these components, along with the algorithms outlining this procedure, are presented in Algorithm 1.
Algorithm 1 Multi Stage Balanced Distillations
2:Output: The fine-tuned student model $M_{s}^{K}$
3:for each stage $k=0,\dots,k-1$ do
[…]
5:     for each domain tail domain $j$ do
6:         Add remaining $x_{j}$ from $D$ to $T$
7:         $\hat{x_{j}}=M_{t}(P_{c},j)$
8:         Add synthetic $\hat{x_{j}}$ to $T$
9:     for each domain head domain $h$ do
10:         Collect all $x_{h}$ from $D$
11:         $x_{h}=M_{s}^{k-1}(x_{h},h)$
12:         Add selected $x_{h}$ to $T$
13:     Use $M_{t}$ to annotate $x$ in $T$ w/o rationales
14:     $M_{s}^{k}=\text{Fine-tune}(M_{s},T)$
@@ 3.5 Student Active Selection
At last, we rank the inputs by their IFD scores, selecting those with the highest values to include in the batch, as specified by the balancing policy.
@@ 3.6 Reasoning Generation and Fine-tuning
After gathering the required samples and their associated rationales in the training batch, we integrate this batch with the annotated data accumulated from previous stages. This approach ensures that our student model is exposed to a diverse and comprehensive dataset, which helps mitigate the risk of overfitting — a common challenge in machine learning models as identified in prior studies Dor et al. (2020); Liu et al. (2023b). To facilitate this, we reinitialize and fine-tune the student model on the compiled rationale sequences from scratch at each stage.
@@ 4 Experiment
To verify the effectiveness of our framework on various reasoning tasks, we evaluate our method on five long-tailed datasets, following previous work Yu et al. (2023); Dai et al. (2023); Huang et al. (2021). These include text classification: R52 and Reuters Hayes and Weinstein (1990), question answering: AbstractiveQA and Multiple-choiceQA Dai et al. (2023) and arithmetic: MATH Hendrycks et al. (2021). For text classification datasets, we treat the label of inputs as the domain; for other datasets, the domain information of inputs is annotated as metadata from the data provider. The detailed construction process and domain information for these datasets can be found in Appendix A. We also show two example distributions of the datasets in Figure 5 in Appendix A. For each dataset, we prepare two budget settings for the experiment. In Table 1, we present the budget number, the test number, the domain number, and the evaluation metric of all five datasets. The budget number in Table 1 represents the total number of queries to the teacher models. For example, budget setting 1 for the R52 dataset is 2,600, which means that for our BalDistill method, the sum of queries to the teacher model for data augmentation and for generating reasoning steps should also equal 2,600 without incurring additional costs. This budget ensures that all operations, including data augmentation and reasoning step generation, are performed within the allocated query limit.
@@
Since we are dealing with long-tailed imbalanced data, for each dataset, we choose to use both the micro- and macro-averages to evaluate the method robustness Henning et al. (2022); Li et al. (2024d). For the classification datasets (R52 and Reuters), we report micro-F1 and macro-F1, where micro-F1 is a global average F1 score and macro-F1 is computed by taking the unweighted mean of all the per-class F1 scores Harbecke et al. (2022). For other datasets, we also report the micro-/macro-F1 for AbstractiveQA datasets and micro-/macro-accuracy for Math and Multi-choiceQA datasets. Note that the F1 score for the AbstractQA is the word-level F1 score between the token list of ground truth answer and the generated answer, different from the F1 for the classification task.
@@
We experiment with two variants of our proposed method with different balancing policies, as discussed in Section 3: In our first framework BalDistill (N), we use naive balancing policy, and for second framework BalDistill (A), we leverage adaptive balancing. We compare our framework with multiple baseline methods: (1) Zero-shot CoT. We directly prompt the student model to infer on the test data Kojima et al. (2022). (2) Random Finetune. We randomly collect samples from the training data until the budget constraint is met and finetune student models on the final ground-truth labels Radford et al. (2019). (3) Random Finetune-CoT. We randomly collect and use CoT rationales from the teacher model for student fine-tuning Ho et al. (2022); Yao et al. (2022); He et al. (2023). (4) Duplicate Finetune-CoT. We construct the training data with a naive balancing policy. For the tail domains, we duplicate the inputs to satisfy the policy requirement and for head domains, we randomly sample examples in over-represented domains.
@@ 5.1 Comparison with Baseline Methods
To perform a detailed analysis of our framework, we visualize the F1 or accuracy score for each domain of the BalDistill (N) method and two baseline methods (Random Finetune CoT and Duplicate Finetune CoT) in Figure 2, with the x-axis representing the proportion of each domain in the dataset in descending order. From Figure 2, our proposed method can achieve comparable results in the head domains (left side of the figure) but substantially outperform the baseline methods in the tail domains (right side of the figure). This observation verifies our expectation in Section 3, where the balancing policy increases performance in the tail domain, and the active learning part improves the data efficiency to compensate for data loss in the head domain. Note that for Math dataset, BalDistill can only achieve comparable results with the baseline methods on the last two tail domains (precalculus and probability), and we conjecture that the high difficulty in these two domains prevents the teacher from composing high-quality synthetic data. We also include 2 additional SoTA methods for multitask learning or resolving class imbalance challenges and compare their results with the performance of BalDistill. Our approach outperforms the baselines in most tasks. Details are shown in Appendix C.
@@ 5.2 Ablation Study
After showing the superiority of our overall framework, our next step is to verify the effectiveness of each component in the proposed method. We compare our framework with the ablated methods: (1) Balance Finetune CoT. We adopt a naive balancing policy to construct the training set and query the teacher model to compose inputs in the tail domains. We randomly sample examples from head domains to make sure they are not over-represented. (2) Active Finetune CoT. We only keep the active learning component but remove the data augmentation part. In details, we calculate the IFD scores for all examples in our original dataset and select the highest IFD scores to satisfy the budget number. Note that this ablation method is equivalent to the SoTA active learning method: Superfiltering Li et al. (2024a). The experiment setting is similar to the setup in Section 5.1, and we present the performance of each method with two budget settings in Table 3.
@@ 5.3.2 Sensitivity Analysis
We next investigate how the choice of stage number: $K$ will influence the performance of our framework. We experiment with the same setup as in Section 5.1 but with varying stage numbers among $\{3,5,8\}$. We visualize the results (macro-averages) of BalDistill (A) and the baseline method Random Finetune CoT in Figure 4
@@
(c) MultiChoiceQA
(d) Math
Figure 4: Influence of stage number choices on BalDistill across datasets. Our proposed method consistently obtains better results than the random fine-tune baseline method with varying stage numbers.
@@ 6 Conclusions
Acknowledgments
Zhou, Wang, Liu and Huang are supported by DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) 80321, National Science Foundation NSF-IIS-2147276 FAI, DOD-ONR-Office of Naval Research under award number N00014-22-1-2335, DOD-AFOSR-Air Force Office of Scientific Research under award number FA9550-23-1-0048, DOD-DARPA-Defense Advanced Research Projects Agency Guaranteeing AI Robustness against Deception (GARD) HR00112020007, Adobe, Capital One and JP Morgan faculty fellowships. Zhu and Koutra are supported by the National Science Foundation CAREER Grant No. IIS 1845491. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or other funding parties.
Limitations
@@ Acknowledgments
Another future direction for our paper is to explore the application of knowledge distillation in Large Vision-Language Models (LVLMs) Liu et al. (2024c, b); Bai et al. (2023); Sun et al. (2023b); Zhou et al. (2024a); Wang et al. (2024a); Lin et al. (2024); Zhu et al. (2024); Liu et al. (2024e). In this paper, we have focused on experiments related to knowledge distillation in Large Language Models (LLMs). In future work, we aim to use knowledge distillation to further address the issue of hallucination Liu et al. (2023a); Cui et al. (2023); Wang et al. (2024b) in small LVLMs such as LLaVA-7b Liu et al. (2024c) and VILA-7b Lin et al. (2024).
@@ Appendix B Implementation Details
We use the zero-shot prompts for the teacher to give the rationales and the few-shot ICL to generate the synthetic tail samples. The prompts are shown in Tables 7, 8 and 9. We call the gpt-4 function from OpenAI to obtain teacher responses.
@@
Appendix C Additional Baseline Results
Besides the baselines mentioned in Section 5.1, we also include two additional state-of-the-art methods for multitask learning and addressing class imbalance challenges. The first method is Glee Zhang et al. (2022), which leverages prompt tuning on masked tokens to handle long-tailed classification tasks. We adapted the classification head of Glee to make it applicable for generation tasks. However, since the final answers for AbstractiveQA consist of multiple words with variable lengths, Glee cannot be leveraged for AbstractiveQA. The second method is MFTCoder Liu et al. (2024a), which proposes various loss functions to address challenges in multitask learning. For our experiment, we treat each domain as a separate task within their framework. Note that we employ instruction tuning for both MFTCoder and BalDistill, while using prompt tuning for Glee.
Below we present the macro metrics of the baselines and BalDistill on each dataset. Llama2 is used as backbones for all of the methods.
Dataset R52 Reuters AbsQA MCQA Math
MFTCoder 10.49 6.63 33.76 58.62 3.22
Glee 46.87 23.96 N/A 57.88 4.37
BalDistill (A) 58.33 25.51 47.55 59.14 4.21
Table 6: Performance of BalDistill and two additional baselines across 5 datasets.
From the results in Table 6, it is evident that our approach outperforms the MFTcoder and Glee on most tasks. MFTCoder’s underperformance can be attributed to its use of validation loss gradients (as described in Equation 4 of their paper) to adjust training loss, which leads to unstable learning. In our experimental setup, particularly for tail labels in the R52 and Reuters datasets, we often had only one or two examples in the validation set. This scarcity can cause large gradients in the validation data, potentially leading to loss explosion during fine-tuning on these datasets. Glee’s limitations results in its failure to utilize information from teacher rationales. In contrast, our method leverages data augmentation from teachers to elicit more knowledge and enhance fine-tuning for tail domains. This approach allows us to better capture and utilize the expertise embedded in the teacher models, resulting in improved performance across various tasks.